# On Joint Noise Scaling in Differentially Private Federated Learning with Multiple Local Steps

### v1

@@ 3 Background
where the sum inside $TA$ is done by a trusted aggregator, $clip_{C}$ ensures that each client-specific update has bounded $\ell_{2}$-norm, and $\xi^{(t)}_{i}$ is Gaussian noise s.t. $\sum_{i=1}^{N}\xi^{(t)}_{i}$ gives the joint DP protection level we are aiming for. As the clipping and noise are applied directly to the updated weights after the local optimisation has finished, the privacy protection is not affected by the number of local optimisation steps client $i$ is using to arrive at $\Delta_{i}^{(t)}$ before applying DP.
There is also a simple approach that works with more fine-grained granularities, when the clients use a single local optimisation step with common learning rate $\gamma$ and, for example, standard DP stochastic gradient descent (DP-SGD, [40]) again utilising Gaussian noise: we can take $\Delta_{i}^{(t)}=-\gamma(g_{i}^{(t)}+\xi_{i}^{(t)})$, where $g_{i}^{(t)}$ is a sum of clipped per-unit gradients (e.g. per-sample for sample-level DP) from client $i$, to have the update
@@
Looking at the sum in Equation 6, since each per-unit gradient has a common bounded norm and Gaussian noise is infinitely divisible, i.e., the summed-up noise is another Gaussian, we can calculate the resulting privacy with standard techniques (see, e.g., [34, 27, 53]).
If one tries to use the same reasoning as with a single step, e.g., with sample-level DP using $S>1$ local optimisation steps, the problem is that the sensitivity of the per-sample clipped gradients when summed over the local steps increases with $S$: assuming $\|g_{i,s}\|_{2}\leq C,s=1,\dots,S$ implies that $\|\sum_{s=1}^{S}g_{i,s}\|_{2}\leq SC$ (triangle-inequality); in other words, trying to scale the noise over the local steps naively ends up scaling the query sensitivity linearly with the total number of steps.
As noted before, the obvious problem in using only a single local optimisation step per FL round is that the number of communication rounds is often one of the main bottlenecks in FL [25]. Furthermore, under some assumptions it is also possible to theoretically show that increasing the number of local steps leads to faster convergence (see [29, Theorem 3.2]). Our novel analysis therefore enables provably faster convergence under some assumptions (e.g., when the assumptions of [29, Theorem 3.2] as well as those stated in Assumption 4.1 hold). More practically, we also empirically demonstrate that taking more local steps can lead to significant gains under various settings in Section 5.
@@ 3.1 Trust Model
In this paper, we assume an honest-but-curious (hbc) server and that all the clients are fully honest. The latter assumption can be easily generalised to allow for hbc clients with some weakening to the relevant privacy bounds: with $N$ (non-colluding) hbc clients, since any client could potentially remove its own noise from the aggregated results, the noise from the other $N-1$ clients needs to guarantee the target DP level. In effect, to allow for all hbc clients, we would need to scale up the noise level somewhat (see, e.g., [22] for a discussion on noise scaling and for formal proofs).
@@
4 Joint Noise Calibration with Multiple Local Steps Using a Trusted Aggregator
Consider standard FL setting with $M$ clients and client $i$ holding some local data $x_{i}$. On FL round $t$, $N_{t}$ clients are selected for updating by the server, w.l.o.g. assumed to be clients $i=1,\dots,N_{t}$. Each selected client $i$ receives the current model parameters $\theta^{(t-1)}$ from the server, then runs $S_{t}$ local optimisation steps using DP-SGD with constant learning rate $\gamma_{t}$, and finally sends an update to the server via a trusted aggregator $TA$:
$\Delta_{i}^{(t)}$ $=\theta_{i}^{(t)}-\theta^{(t-1)}$ (7)
$=-\sum_{s=1}^{S_{t}}\gamma_{t}(g^{(t)}_{i,s}+\xi^{(t)}_{i,s}),$ (8)
@@ 4 Joint Noise Calibration with Multiple Local Steps Using a Trusted Aggregator
In the rest of this section we state our main results: we show that under some assumptions we can account for privacy in FL by looking at the local optimisation steps while scaling the noise level jointly over the clients, even if there is no communication between the clients during the local optimisation but only a single trusted aggregation at the end of the round to update the global model parameters.
W.l.o.g. from now on we drop the FL round index $t$ and simply write, e.g., $N$ instead of $N_{t}$ for the number of updating clients. Since the global updates do not access any sensitive data, once we can do privacy accounting for a single FL round, which is the main topic in the rest of this section, generalising to $T$ FL rounds can be done in a straightforward manner. See Appendix Appendix for more details on the privacy accounting we use.
In the following, we assume that all clients have access to an ideal trusted aggregator, and that all sums are calculated by calling the trusted aggregator. We comment on more realistic implementations after stating our main results.
@@
Let $x_{i}\in\mathcal{X}^{*},i=1,\dots,N$. We write $x=\cup_{i=1}^{N}x_{i}$, and assume that $x_{i}\cap x_{j}=\varnothing$ for every $i\not=j$, i.e., there are no overlapping samples in different clients’ local data sets. We are interested in fixed-length optimisation runs of $S$ local steps (common to all clients), which leads to (fixed-length) adaptive sequential composition for privacy accounting (see e.g. [36, 53]). We assume all clients use the same learning rate $\gamma$ and norm clipping with constant $C$ when applicable. We also assume that all local DP mechanisms ${\mathcal{A}}^{(s)}_{i},s=1,\dots,S,i=1,\dots,N$ are DP w.r.t. the first argument for any given auxiliary values (which we generally do not write out explicitly).
Note that we consider how to loosen many of these assumptions in Appendix Appendix.
Not all possible DP mechanisms might allow for joint noise scaling via simple aggregation. For convenience, in Definition 4.2 we define a family of suitable mechanisms, which we call sum-dominating mechanisms:
@@
Let ${\mathcal{A}},{\mathcal{A}}_{i}:\mathcal{X}^{*}\rightarrow\mathcal{O},i=1,\dots,N$ be randomised algorithms that satisfy DP. We call ${\mathcal{A}}$ a sum-dominating mechanism, if
$\sup_{x\simeq x^{\prime}}H_{\alpha}\left(\sum_{i=1}^{N}{\mathcal{A}}_{i}(x_{i})\|\sum_{i=1}^{N}{\mathcal{A}}_{i}(x^{\prime}_{i})\right)$ (10)
$\leq$ $\sup_{x\simeq x^{\prime}}H_{\alpha}\left({\mathcal{A}}(x)\|{\mathcal{A}}(x^{\prime})\right),$ (11)
where $H_{\alpha}$ is the hockey-stick divergence, and $\simeq$ is the DP neighbourhood relation.
Considering concrete mechanisms that satisfy Definition 4.2, one simple example is given by DP mechanisms that use infinitely divisible noise, as formalised next in Lemma 4.3:
Lemma 4.3 (Additive mechanisms with infinitely divisible noise are sum-dominated).
Assume ${\mathcal{A}}_{i},i=1,\dots,N$ are additive DP mechanisms s.t. they add noise from an infinitely divisible noise family $\Xi$:
${\mathcal{A}}_{i}(x_{i})=f(x_{i})+\xi_{i},$ (12)
where $f$ is some (bounded sensitivity) function, and $\xi_{i}\in\Xi\ \forall i$. Then the mechanism
${\mathcal{A}}(x):=\sum_{i=1}^{N}\left(f(x_{i})+\xi_{i}\right)$ (13)
is a sum-dominating mechanism w.r.t. ${\mathcal{A}}_{i},i=1,\dots,N$.
Proof.
Immediately clear by definition of ${\mathcal{A}}$. ∎
One prominent example of the possible mechanisms covered by Lemma 4.3 is the ubiquitous continuous Gaussian mechanism:
@@
Assume ${\mathcal{A}}_{i}$ is a Gaussian mechanism with noise covariance $C^{2}\sigma_{i}^{2}\cdot I_{d}$ and $f$ has bounded sensitivity $C$. Since the normal distribution is infinitely divisible, from Lemma 4.3 it follows that the combined mechanism ${\mathcal{A}}=\sum_{i=1}^{N}{\mathcal{A}}_{i}$, which is another Gaussian with sensitivity $C$ and noise covariance $C^{2}(\sum_{i=1}^{N}\sigma^{2}_{i})\cdot I_{d}$, is a sum-dominating mechanism. Finally, due to well-known existing results (see e.g. [32, 27, 53]), a (tightly) dominating pair of distributions $(P,Q)$ in the sense of Definition 3.3 for the sum-dominating mechanism ${\mathcal{A}}$ is given by a pair of 1d Gaussians with means $\mu_{P}=0,\mu_{Q}=1$, and variances $\sigma^{2}_{P}=\sigma^{2}_{Q}=\sum_{i=1}^{N}\sigma^{2}_{i}$.
Other mechanisms covered by Lemma 4.3 include existing discrete infinitely divisible noise mechanisms compatible with practical SecAgg protocols, such as Skellam [46, 1], and Poisson-binomial [14]. [^3] Discrete Gaussian [11] is not infinitely divisible, but is close-enough that a sum-dominating mechanism can still be found in many practical settings, see [24]. In such cases, the inequality in Definition 4.2 could always be strict, whereas for any infinitely divisible noise mechanism it can be written as an equality (see Lemma 4.3). We note that even in the infinitely divisible case, however, writing Definition 4.2 with inequality is necessary to avoid nonsensical limitations, such as having a sum-dominating mechanism satisfy $(\varepsilon,\delta)$-DP, while not satisfying $(\varepsilon,\delta^{\prime})$-DP for any $\delta^{\prime}>\delta$.
Next, we consider composing a sum-dominating mechanisms over $S$ (local) steps. This allows us to account for the total privacy when accessing the sensitive data multiple times:
@@
Assume ${\mathcal{A}}^{(s)}$ is a sum-dominating mechanism for ${\mathcal{A}}_{i}^{(s)},i=1,\dots,N$ for every $s=1,\dots,S$. Then the composition of the sum-dominating mechanisms $({\mathcal{A}}^{(1)},\dots,{\mathcal{A}}^{(S)})$ dominates the composition
@@
$\sup_{x\simeq x^{\prime}}H_{\alpha}\left(\sum_{i=1}^{N}{\mathcal{A}}^{(s)}_{i}(x_{i})\|\sum_{i=1}^{N}{\mathcal{A}}^{(s)}_{i}(x^{\prime}_{i})\right)$ (15)
$\leq$ $\sup_{x\simeq x^{\prime}}H_{\alpha}\left({\mathcal{A}}^{(s)}(x)\|{\mathcal{A}}^{(s)}(x^{\prime})\right)$ (16)
by definition of ${\mathcal{A}}$ (Definition 4.2). The claim therefore follows immediately from [53, Theorem 10]. ∎
Considering Lemma 4.5, in our case it essentially says that to account for running $S$ local optimisation steps, it is enough to find a proper sum-dominating mechanism for each step separately.
With the next result given in Lemma 4.6, we can connect the previous results with the form of output we get from actually running local optimisation in FL:
@@
Due to the post-processing immunity of DP (see, e.g., [18]), the assumption implies that releasing
@@
Taken together, Lemmas 4.3, 4.5 & 4.6 allow us to compose DP mechanisms with joint noise scaling over the clients. In our main result given as Theorem 4.7, we show that each client can run DP-SGD with several local steps while still benefiting from joint noise scaling when communicating the update via a trusted aggregator.
@@
Assume $N$ clients use local noise mechanisms ${\mathcal{A}}_{i}^{(s)},i=1,\dots,N$ as in Lemma 4.3 for each local gradient optimisation step $s=1,\dots,S$, and that the final aggregated update $\sum_{i=1}^{N}\Delta_{i}$ is released via an ideal trusted aggregator. Then denoting the sum-dominating mechanism for step $s$ by ${\mathcal{A}}^{(s)}$, the query release satisfies $(\varepsilon(\delta),\delta)$-DP for any $\delta\in[0,1]$, when $\varepsilon(\delta)$ is given by accounting for releasing the vector
@@
From Lemma 4.6 it follows that valid DP guarantees can be established by accounting for the release of the vector $\left(\sum_{i=1}^{N}{\mathcal{A}}_{i}^{(1)}(x_{i}),\dots,\sum_{i=1}^{N}{\mathcal{A}}_{i}^{(S)}(x_{i})\right)$. Furthermore, Lemma 4.3 implies that for any step $s\in{1,\dots,S}$, the sum-dominating mechanism ${\mathcal{A}}^{(s)}$ dominates $\sum_{i=1}^{N}{\mathcal{A}}_{i}^{(s)}$, and therefore by Lemma 4.5 the claim follows. ∎
We note that while Theorem 4.7 assumes infinitely divisible noise mechanism (as is commonly used with DP-SGD in practice), the result is trivial to generalize to any sum-dominating mechanism ${\mathcal{A}}$, such as discrete Gaussian [11], by relying on Definition 4.2 instead of using Lemma 4.3.
Considering tightness of the privacy accounting done based on Theorem 4.7, it is worth noting that since the accounting relies on Lemma 4.6, which assumes releasing each local step while the actually released query answer is a sum over the local steps, the resulting privacy bound need not be tight, but an upper bound on the privacy budget. However, this matches the usual DP-SGD privacy accounting analysis (see e.g. [34, 27]), which typically needs to account for each local optimisation step due to technical reasons even if only the final model is released. In the general case, it has been shown that hiding the intermediate steps does not bring any privacy benefits compared to the per-step accounting [3].
In case all the clients use data subsampling with suitable and matching sampling probabilities during local training, standard privacy amplification results imply improved privacy bounds (see, for example, [41] and the references therein).
For implementing the trusted aggregator in Theorem 4.7 in practice, it should be noted that as the sum over $s$ is done locally by each client during local optimisation, it is always trusted as long as the individual clients are, while the sum over $i$ would need to be implemented, e.g., using a suitable SecAgg protocol, such as [7, 10, 37, 39]. This places some extra requirements on the DP mechanisms ${\mathcal{A}}_{i}^{(s)}$, since SecAgg protocols typically run on elements of finite rings, which precludes continuous noise mechanisms (at least if the goal is to have joint noise scaling via the summation).
A viable alternative is to use some suitable discrete noise mechanism, such as Skellam [1] or Poisson-binomial [14]. However, since the clients send model updates instead of single gradients, the finite ring size used in the SecAgg needs to accommodate the model update size: it does no good to use Skellam mechanism with gradient quantization to 8 bits, if the model weights and the resulting model update $\Delta_{i}$ for client $i$ still uses 32 bit floats.

### v2

@@ 3.2 Problem with Local Steps in DPFL with SecAgg
where the sum inside $TA$ is done by a trusted aggregator, $clip_{C}$ ensures that each client-specific update has $\ell_{2}$-norm bounded by the constant $C$, and $\xi^{(t)}_{i}$ is Gaussian noise s.t. $\sum_{i=1}^{N}\xi^{(t)}_{i}$ gives the DDP protection level we are aiming for. As the clipping and noise are applied directly to the updated weights after the local optimization has finished, the privacy protection is not affected by the number of local optimization steps client $i$ is using to arrive at $\Delta_{i}^{(t)}$ before applying DP.
There is also a simple approach that works with more fine-grained granularities, when the clients use a single local optimization step with common learning rate $\gamma$ and, for example, standard DP stochastic gradient descent (DP-SGD, Song et al. 2013) again utilising Gaussian noise: we can take $\Delta_{i}^{(t)}=-\gamma(g_{i}^{(t)}+\xi_{i}^{(t)})$, where $g_{i}^{(t)}$ is a sum of clipped per-unit gradients (e.g. per-sample for sample-level DP) from client $i$, to have the update
@@
Looking at the sum in Equation 5, since each per-unit gradient has a common bounded norm and Gaussian noise is infinitely divisible, i.e., the summed-up noise is another Gaussian, we can calculate the resulting privacy with standard techniques (see, e.g., Mironov et al. 2019; Koskela et al. 2020; Zhu et al. 2022). Now, if one tries to use the same reasoning with sample-level DP using $S>1$ local optimization steps, the problem is that the sensitivity of the per-sample clipped gradients when summed over the local steps increases with $S$: assuming $\|g_{i,s}\|_{2}\leq C,s=1,\dots,S$ implies that $\|\sum_{s=1}^{S}g_{i,s}\|_{2}\leq SC$ (triangle-inequality). [^4] Note that, as already mentioned in Section 2, for the special case of Gaussian noise with a trusted server, the analysis of Noble et al. (2022) shows that we can indeed avoid the problem and get valid DDP guarantees that improve on the LDP case.
In other words, trying to scale the noise over multiple local optimization steps naively ends up scaling the query sensitivity linearly with the total number of steps, while the obvious problem in using only a single step per FL round is that the number of communication rounds is typically one of the main bottlenecks in FL (Kairouz et al., 2019).
@@ 3.3 Trust Model
In this paper, we assume an honest-but-curious (hbc) server and that all the clients are fully honest. The latter assumption can be easily generalised to allow for hbc clients with some weakening to the relevant privacy bounds: with $N$ (non-colluding) hbc clients, since any client could potentially remove its own noise from the aggregated results, the noise from the other $N-1$ clients needs to guarantee the target DP level. In effect, to allow for all hbc clients, we would need to scale up the noise level somewhat (see, e.g., Heikkilä et al. 2017 for a discussion on the noise scaling and for formal proofs).
@@
4 Privacy Accounting for Multiple Local Steps Using a Trusted Aggregator
Consider standard FL setting with $M$ clients and client $i$ holding some local data $x_{i}$. On FL round $t$, $N_{t}$ clients are selected for updating by the server, w.l.o.g. assumed to be clients $i=1,\dots,N_{t}$. Each selected client $i$ receives the current model parameters $\theta^{(t-1)}$ from the server, then runs $S_{t}$ local optimization steps using DP-SGD with constant learning rate $\gamma_{t}$, and finally sends an update to the server via a trusted aggregator $TA$:
$\Delta_{i}^{(t)}=\theta_{i}^{(t)}-\theta^{(t-1)}=-\sum_{s=1}^{S_{t}}\gamma_{t}(g^{(t)}_{i,s}+\xi^{(t)}_{i,s}),$ (6)
@@ 4 Privacy Accounting for Multiple Local Steps Using a Trusted Aggregator
In the rest of this section we state our main results: we show that under some assumptions we can account for privacy in FL by looking at the local optimization steps while benefiting from the noise added by all the clients on each step, even if there is no communication between the clients during the local optimization but only a single trusted aggregation at the end of the round to update the global model parameters.
W.l.o.g. from now on we drop the FL round index $t$ and simply write, e.g., $N$ instead of $N_{t}$ for the number of updating clients. Since the global updates do not access any sensitive data, once we can do privacy accounting for a single FL round, which is the main topic in the rest of this section, generalising to $T$ FL rounds can be done in a straightforward manner (see Appendix A.5).
In the following, we assume that all clients have access to an ideal trusted aggregator, and that all sums are calculated by calling the trusted aggregator. We comment on more realistic implementations in Appendix A.4 after stating our main results.
@@
Let $x_{i}\in\mathcal{X}^{*},i=1,\dots,N$. We write $x=\cup_{i=1}^{N}x_{i}$, and assume that $x_{i}\cap x_{j}=\varnothing$ for every $i\not=j$, i.e., there are no overlapping samples in different clients’ local data sets. We are interested in fixed-length optimization runs of $S$ local steps (common to all clients), which leads to (fixed-length) adaptive sequential composition for privacy accounting (see e.g. Rogers et al. 2016; Zhu et al. 2022). We assume all clients use the same learning rate $\gamma$ and norm clipping with constant $C$ when applicable. We also assume that all local DP mechanisms ${\mathcal{A}}^{(s)}_{i},s=1,\dots,S,i=1,\dots,N$ are DP w.r.t. the first argument for any given auxiliary values (which we generally do not write out explicitly).
Note that we consider how to loosen many of these assumptions in Appendix A.3.
Not all possible DP mechanisms might allow for improved DDP guarantees via simple aggregation. For convenience, in Definition 4.2 we define a family of suitable mechanisms, which we call sum-dominating:
@@
Let ${\mathcal{A}},{\mathcal{A}}_{i}:\mathcal{X}^{*}\rightarrow\mathcal{O},i=1,\dots,N$ be randomised algorithms. We call ${\mathcal{A}}$ a sum-dominating mechanism w.r.t. ${\mathcal{A}}_{i},i=1,\dots,N$, if
$\sup_{x\simeq x^{\prime}}H_{\alpha}\left(\sum_{i=1}^{N}{\mathcal{A}}_{i}(x_{i})\|\sum_{i=1}^{N}{\mathcal{A}}_{i}(x^{\prime}_{i})\right)\leq\sup_{x\simeq x^{\prime}}H_{\alpha}\left({\mathcal{A}}(x)\|{\mathcal{A}}(x^{\prime})\right),$ (8)
where $H_{\alpha}$ is the hockey-stick divergence, and $\simeq$ is the DP neighbourhood relation. If additionally $\sup_{x\simeq x^{\prime}}H_{\alpha}\left({\mathcal{A}}(x)\|{\mathcal{A}}(x^{\prime})\right)<\max_{i}\sup_{x\simeq x^{\prime}}H_{\alpha}\left({\mathcal{A}}_{i}(x_{i})\|{\mathcal{A}}_{i}(x_{i}^{\prime})\right)$, then we call ${\mathcal{A}}$ a properly sum-dominating mechanism.
Looking at Definition 4.2, in practice the interesting mechanisms for us will be the properly sum-dominating mechanisms, as these have improved privacy bounds compared to any single mechanism ${\mathcal{A}}_{i}$ in isolation. However, from the following results only Theorem 4.7 is specific to proper sum-dominating mechanisms. Considering concrete mechanisms that are properly sum-dominating, simple examples are given by the existing DP mechanisms that use infinitely divisible noise (see Definition A.1), such as Skellam (Valovich & Aldà, 2017; Agarwal et al., 2021), Poisson-binomial (Chen et al., 2022b), as well as the ubiquitous continuous Gaussian (Dwork et al., 2006a): [^5] Discrete Gaussian (Canonne et al., 2020) is not infinitely divisible, but is close-enough that a properly sum-dominating mechanism can still be found in many practical settings, see Kairouz et al. (2021). In such cases, the inequality in Equation 8 could always be strict, whereas for the infinitely divisible noise mechanisms mentioned in the text it can be written as an equality. We note that even in this case, however, writing Equation 8 with inequality is necessary to avoid nonsensical limitations, such as having a DP mechanism that satisfies Definition 4.2 with a given $\delta$ while not satisfying it for any $\delta^{\prime}>\delta$.
@@
Assume ${\mathcal{A}}_{i}$ is a Gaussian mechanism with noise covariance $C^{2}\sigma_{i}^{2}\cdot I_{d}$ and $f$ has bounded sensitivity $C$. Since the normal distribution is infinitely divisible (Definition A.1), the combined mechanism ${\mathcal{A}}=\sum_{i=1}^{N}{\mathcal{A}}_{i}$ is another Gaussian with sensitivity $C$ and noise covariance $C^{2}(\sum_{i=1}^{N}\sigma^{2}_{i})\cdot I_{d}$. Finally, due to well-known existing results (see e.g. Meiser & Mohammadi 2018; Koskela et al. 2020; Zhu et al. 2022), a (tightly) dominating pair of distributions $(P,Q)$ in the sense of Definition 3.3 for the sum-dominating mechanism ${\mathcal{A}}$ is given by a pair of 1d Gaussians with means $\mu_{P}=0,\mu_{Q}=1$, and variances $\sigma^{2}_{P}=\sigma^{2}_{Q}=\sum_{i=1}^{N}\sigma^{2}_{i}$. As the DP guarantees for the Gaussian improve when increasing the variance (Balle & Wang, 2018), ${\mathcal{A}}$ is a properly sum-dominating mechanism w.r.t. to ${\mathcal{A}}_{i},i=1,\dots,N$ (Definition 4.2).
Next, we consider composing a sum-dominating mechanisms over $S$ local steps. This allows us to account for the total privacy when doing more than one local optimization step:
@@
Assume ${\mathcal{A}}^{(s)}$ is a sum-dominating mechanism w.r.t. ${\mathcal{A}}_{i}^{(s)},i=1,\dots,N$ for every $s=1,\dots,S$. Then the composition of the sum-dominating mechanisms $({\mathcal{A}}^{(1)},\dots,{\mathcal{A}}^{(S)})$ dominates the composition
@@
$\sup_{x\simeq x^{\prime}}H_{\alpha}\left(\sum_{i=1}^{N}{\mathcal{A}}^{(s)}_{i}(x_{i})\|\sum_{i=1}^{N}{\mathcal{A}}^{(s)}_{i}(x^{\prime}_{i})\right)\leq\sup_{x\simeq x^{\prime}}H_{\alpha}\left({\mathcal{A}}^{(s)}(x)\|{\mathcal{A}}^{(s)}(x^{\prime})\right)$ (10)
by definition of ${\mathcal{A}}$ (Definition 4.2). The claim therefore follows immediately from (Zhu et al., 2022, Theorem 10). ∎
Considering Lemma 4.4, in our case it essentially says that to account for running $S$ local optimization steps, it is enough to find a sum-dominating mechanism for each step separately.
With the next result given in Lemma 4.5, we can connect the previous results with the form of output we get from actually running local optimization in FL:
@@
Due to the post-processing immunity of DP (see, e.g., Dwork & Roth 2014), the assumption implies that releasing
@@
Taken together, Definition 4.2 along with Lemmas 4.4 & 4.5 allow us to compose DP mechanisms for DDP guarantees using LDP noise from multiple clients. In our main result given as Theorem 4.6, we show that with (properly) sum-dominating mechanisms each client can run DP-SGD with several local steps and LDP noise while receiving (improved) DDP guarantees when communicating the update via a trusted aggregator.
@@
Assume $N$ clients use local noise mechanisms ${\mathcal{A}}_{i}^{(s)},i=1,\dots,N$ as in Definition 4.2 for each local gradient optimization step $s=1,\dots,S$, and that the final aggregated update $\sum_{i=1}^{N}\Delta_{i}$ is released via an ideal trusted aggregator. Then denoting the sum-dominating mechanism for step $s$ by ${\mathcal{A}}^{(s)}$, the query release satisfies $(\varepsilon(\delta),\delta)$-DP for any $\delta\in[0,1]$, when $\varepsilon(\delta)$ is given by accounting for releasing the vector
@@
From Lemma 4.5 it follows that valid DP guarantees can be established by accounting for the release of the vector $\left(\sum_{i=1}^{N}{\mathcal{A}}_{i}^{(1)}(x_{i}),\dots,\sum_{i=1}^{N}{\mathcal{A}}_{i}^{(S)}(x_{i})\right)$. Furthermore, Definition 4.2 implies that for any step $s\in{1,\dots,S}$, the sum-dominating mechanism ${\mathcal{A}}^{(s)}$ dominates $\sum_{i=1}^{N}{\mathcal{A}}_{i}^{(s)}$, and therefore by Lemma 4.4 the claim follows. ∎
Considering properly sum-dominating mechanisms based on infinitely divisible noise, we also have the following theorem on scaling the noise with multiple clients:
Theorem 4.7.
In addition to the assumptions of Theorem 4.6, assume that each client uses a common noise level $\sigma_{LDP}$ for some mechanism based on infinitely divisible additive noise for each local step s.t. the DP noise $\xi_{LDP}$ for any given local mechanism ${\mathcal{A}}_{i}^{(s)},i=1,\dots,N,s=1,\dots,S$, has covariance $\sigma_{LDP}^{2}\Sigma_{d}$ for some $\Sigma_{d}$. Then for guaranteeing given $(\varepsilon(\delta),\delta)$-DDP, the required $\sigma_{LDP}$ scales as ${\mathcal{O}}(\frac{1}{\sqrt{N}})$.
Proof.
Let $\sigma^{2}\Sigma$ be such that when accounting for the total of $S$ local steps, the composition of the sum-dominating mechanisms $\left({\mathcal{A}}^{(1)},\dots,{\mathcal{A}}^{(S)}\right)$ is $(\varepsilon(\delta),\delta)$-DDP. From Definition A.1 it immediately follows that the required total noise can be broken into parts while staying in the same noise family to give
$\sigma^{2}=\sum_{i=1}^{N}\sigma_{LDP}^{2}\Leftrightarrow\sigma_{LDP}^{2}=\frac{\sigma^{2}}{N}\Leftrightarrow\sigma_{LDP}=\frac{\sigma}{\sqrt{N}}.$
We note that similar scaling results as given in Theorem 4.7 for specific cases like the Gaussian mechanism are already well-known in the existing literature (see, e.g., Noble et al. 2022; Heikkilä et al. 2017).
Finally, considering tightness of the privacy accounting done based on Theorem 4.6, it is worth noting that since the accounting relies on Lemma 4.5, which assumes releasing each local step while the actually released query answer is a sum over the local steps, the resulting privacy bound need not be tight but an upper bound on the privacy budget. However, this matches the usual DP-SGD privacy accounting analysis (see e.g. Mironov et al. 2019; Koskela et al. 2020), which typically needs to account for each local optimization step due to technical reasons even if only the final model is released. In the general case, it has also been shown that hiding the intermediate steps does not bring any privacy benefits compared to the per-step accounting (Annamalai, 2024).